Caroline's age is 3 times her daughter's. In eight years, the sum of their ages will be 72. How old are they now?

Answer by ultrakimidis1(8)

Daughter's age=x
Caro's age= 3x
In 8 years time, Caro's age =3x+8
Daughter's age= x+8
Sum of both ages are 72
3x+8+x+8=72
4x+16=72
4x=72-16
4x=56
x=14

Daughter's age =14
Caro's age=3(14)= 42